# April 2026

### 04 Apr 2026 v1

@@ Abstract
We study the role of commutativity in the representation of relativistic locality for localization observables of relativistic quantum systems in Minkowski spacetime. A celebrated no-go theorem by Halvorson and Clifton shows that commutativity of localization effects associated with causally separated regions is incompatible with other apparently natural assumptions on spatial localization. Commutativity is in particular assumed to be the mathematical representation of locality in the Araki-Haag-Kastler formulation of QFT. This raises the question of whether commutativity follows from more elementary locality principles of Quantum Theory. Adopting Busch’s operational analysis in terms of no-signaling and relativistic consistency, we argue that, for a particle-like system, the commutativity requirement is not a consequence of these principles. Under a natural local detectability principle, elementary localization observables are not localized in arbitrarily small spacetime neighborhoods of the corresponding spatial regions, but rather in regions containing the whole rest space (a Cauchy surface) where the measurement is performed. This is a consequence of the very nature of particle which is assumed to localize in a unique place on a rest space completely filled with ideal detectors. To this regard there is no direct conflict with Araki-Haag-Kastler formulation of local quantum physics. On the other hand we also shows that commutativity and localization may coexists if referring to less idealized localization procedures. Indeed, we introduce conditional localization POVMs associated with bounded spatial regions interpreted as laboratories. These observables, due to a result in quantum information theory known as gentle measurement lemma, describe conditional localization probabilities. In principle, these observables satisfy commutativity when associated with causally separated laboratories. As a consequence they could be represented by local observables in the sense of Araki-Haag-Kastler. Explicit examples of these local observables satisfying a commutative requirement will be presented is a forthcoming paper within the approach of local QFT.
@@ 1.1 Localization, Causality, Commutativity
In non-relativistic quantum mechanics, localization is naturally described by projection-valued measures and, correspondingly, by selfadjoint position operators. This is a case of sharp localization, since the projectors $P(\Delta)$ of the spectral measure are labelled by arbitrarily small regions $\Delta$ of the rest space $\Sigma$ of an observer, and we can prepare states $\rho$ for which the probability of detecting a particle a given such bounded region is $tr(\rho P(\Delta))=1$.
@@
From the one-particle viewpoint, a further serious issue is raised by Hegerfeldt’s celebrated analysis [Heg74, Heg85]. If energy, i.e., the selfadjoint generator of time translations is positive (or bounded from below), sharp localization described by projection-valued measures gives rise to causality problems. In modern terms, and especially in the formulation later clarified by Castrigiano, these problems concern the incompatibility of sharp localization with a natural requirement of causal (i.e. non-superluminal) propagation for localization probabilities. This strongly suggests that, in relativistic quantum theory, sharp notions of localization should be abandoned in favour of unsharp ones described by POVMs. In this case, effects, that is, positive operators bounded by the identity, $0\leq A(\Delta)\leq I$, rather than orthogonal projectors, are still associated with arbitrarily small regions $\Delta$ in the rest space $\Sigma$ of an observer, but the probability of finding the particle in a given bounded region is strictly $tr(\rho A(\Delta))<1$ for every state $\rho$, even though sharp states can be approximated arbitrarily well.
Recent works [Cas17, Mor23, Cas24, DRMo24, Cas25, CDM26] have shown that this strategy is in fact viable. A broad family of relativistic spatial localization observables has been constructed and analyzed for various elementary systems, with full compatibility with Poincaré covariance and with natural causal requirements, especially in the unsharp setting when the reqirement on positive energy is assumed. These developments indicate that Hegerfeldt’s issue does not rule out unsharp localization under the requirement of energy positivity, but rather rules out an excessively rigid, projection-based notion of localization.
@@
Actually it is not even supposed that a whole unitary representation of the Poincaré group exists. The relativistic perspective comes into the play just to physically justify commutativity of localization effects associated to ”separated” regions, when interpreted as causally separated regions in relativistic sense.
@@
The crucial point is that the commutativity requirement entering the Halvorson-Clifton theorem is usually regarded as self-evident from the perspective of local quantum physics, especially within the Araki-Haag-Kastler framework [Haa96, Ara09], where locality is encoded precisely by commutativity of observables associated with causally separated spacetime regions. If the spatial region $\Delta$ of a restapace $\Sigma$ where a particle is found at a certain time $t$ belongs to the spacetime region $\cal O$} where the localization-observable $A(\Delta)$ is localized (in AHK sense), then the commutativity assumption would indeed appear unavoidable. From this standpoint Halvorson-Clifton’s result would represent an insurmountable obstruction to any notion of position of a quantum system at given time in a reference frame. However, this line of reasoning already presupposes a specific mathematical representation of locality. It is therefore natural to investigate the interplay of commutativity and causally separation of the localization region, independently of the assumption that localization observables $A(\Delta)$ belong to AHK local algebras ${{\cal A}}({\cal O})$, is strict sense, attached to arbitrarily small spacetime neighborhoods $\cal O$ of the corresponding spatial regions: ${\cal O}\supset\Delta$. A first step in this direction was taken by Busch [Bus99], building on a technical result with Singh [BuSi98], by analyzing the relation between commutativity and locality directly within the general theory of quantum measurement. In Busch’s approach an observable is localized in a region $\cal O$} of spacetime if it can be measured with instruments which work therein. This is also in the spirit of AHK structure, but contrarily to that framework, no further requirement about the structure (*, $C^{*}$, von Neumann algebra) is assumed for the family of observables localized in a region of spacetime. Similarly, no a priori requirements for families of observables associated to distinct regions of spacetime are supposed. In that setting, locality is formulated through operational principles such as the no-signaling condition and, in a stronger form, what is often called relativistic consistency [Bec20], namely invariance of the statistics of suitable consecutive measurements under interchange of their temporal order when the measurements take place in causally separated regions. For broad classes of measurement procedures – in particular for Lüders measurements and related efficient schemes – these principles imply commutativity of the relevant effects. This analysis is mathematically precise and conceptually important because it does not assume locality in the algebraic form of microcausality (i.e. commutativity), but rather attempts to derive commutativity from independent operational requirements.
@@
The above negative answer is not due to a failure of no-signaling or relativistic consistency, but rather in the physical meaning of localization for a quantum system which we define to be a particle. Let us assume that a localization observable is meant to describe a system with the propensity to be found in a unique place when an ideal position measurement is performed on a whole rest space. Then the elementary observable associated with detection in a region $\Delta$ is not naturally localized in an arbitrarily small spacetime neighborhood of $\Delta$. On the contrary, once one adopts the above stated local detectability principle, as we shall prove in Section 4.2, one is led to the conclusion that the elementary observable $\{A(\Delta),I-A(\Delta)\}$ is localized, in Busch sense, in a region that must contain the whole rest space where the global localization experiment is performed. As a consequence, two elementary localization observables associated with spatially disjoint regions are not localized in causally separated spacetime regions, and therefore the no-signaling and relativistic-consistency arguments leading to commutativity cannot even be triggered. In this sense, the commutativity requirement for localization effects does not follow from those more elementary locality principles, since the very hypotheses for employing these principles are not satisfied. This conclusion is physically tied to the particle character of the systems under consideration. A particle is assumed not to localize simultaneously in several (causally) disjoint regions in an ideal exhaustive detection experiment, but rather in a unique place. If one measures whether the particle is found in a region $\Delta$ within a complete position measurement on a rest space $\Sigma$, then one automatically obtains information about the complementary region $\Sigma\setminus\Delta$. From this viewpoint, the associated elementary observable is not confined to $\Delta$ alone: it carries information about the global detection arrangement on $\Sigma$. The failure of the commutativity argument is therefore not an exotic mathematical pathology, but rather a direct consequence of the physical meaning attributed to localization of a particle. If we do not accept this, we cannot accept the very notion of a particle.
These arguments provide a motivation, for asserting that localization effects, for particles, do not belong to local algebras in the sense of AHK formulation, in the case where they are parts of a POVM which fills the whole rest space of a reference frame (a Cauchy surface).
@@
We introduce a notion of conditional localization associated with a bounded spatial region $\Delta_{0}$, interpreted as the spatial part of a laboratory. The idea is to construct, starting from a general localization observable defined on a full rest space, a new POVM normalized on $\Delta_{0}$ which describes the conditional probability of finding the system in a subregion $\Delta\subset\Delta_{0}$, given that the system has been detected in $\Delta_{0}$. The mathematical possibility of such construction relies upon a known result in quantum information theory called gentle measurement lemma. It permits to prove that a specific type of POVMs we introduce in a laboratory normalized on bonded spatial region $\Delta_{0}$, measure the conditional probability to find a particle in $\Delta\subset\Delta_{0}$ if we know that the system was detected in $\Delta_{0}$. This is all the more true the closer the probability of finding the particle in $\Delta_{0}$ is to $1$. These conditional POVMs are intrinsically local to the laboratory and, unlike the original localization observables on the whole rest space, are not affected by the same problem that prevents the use of no-signaling or relativistic-consistency arguments. We shall argue that pairs of such conditional POVMs associated with causally separated laboratories may, in principle, satisfy commutativity and belong to local algebras concentrated around the laboratory region in the AHK sense. This suggests that conditional localization may provide a mathematically and physically appropriate notion of local position measurement in relativistic quantum theory. Explicit examples of these local observables satisfying a commutative requirement and constructed in terms of local energies will be presented is a forthcoming paper within the approach of local QFT.

### 07 Apr 2026 v2

@@ Abstract
We study the role of commutativity in the representation of relativistic locality for localization observables of relativistic quantum systems in Minkowski spacetime. A celebrated no-go theorem by Halvorson and Clifton shows that the commutativity of localization effects associated with causally separated regions is incompatible with other apparently natural assumptions about spatial localization. In particular, commutativity is assumed to provide the mathematical representation of locality in the Araki-Haag-Kastler formulation of quantum field theory. This raises the question of whether commutativity follows from more elementary locality principles of quantum theory. Adopting Busch’s operational analysis in terms of no-signaling and relativistic consistency, we argue that, for a particle-like system, the commutativity requirement does not follow from these principles. Under a natural local detectability principle, elementary localization observables are not localized in arbitrarily small spacetime neighborhoods of the corresponding spatial regions, but rather in regions containing the entire rest space (a Cauchy surface) on which the measurement is performed. This is a consequence of the very nature of a particle, which is assumed to be localized at a unique position on a rest space completely filled with ideal detectors. In this respect, there is no direct conflict with the Araki-Haag-Kastler formulation of local quantum physics. On the other hand, we also show that commutativity and localization may coexist when one refers to less idealized localization procedures. Indeed, we introduce conditional localization POVMs associated with bounded spatial regions interpreted as laboratories. Owing to a result in quantum information theory known as the gentle measurement lemma, these observables describe conditional localization probabilities. In principle, these observables satisfy commutativity when associated with causally separated laboratories. As a consequence, they could be represented by local observables in the sense of Araki-Haag-Kastler. Explicit examples of such local observables satisfying the commutativity requirement will be presented in a forthcoming paper within the framework of local quantum field theory.
@@ 1.1 Localization, Causality, Commutativity
In non-relativistic quantum mechanics, localization is naturally described by projection-valued measures and, correspondingly, by selfadjoint position operators. This is a case of sharp localization, since the projectors $P(\Delta)$ of the spectral measure are labelled by arbitrarily small regions $\Delta$ of the rest space $\Sigma$ of an observer, and we can prepare states $\rho$ for which the probability of detecting a particle in a given such bounded region is $tr(\rho P(\Delta))=1$.
@@
From the one-particle viewpoint, a further serious issue is raised by Hegerfeldt’s celebrated analysis [Heg74, Heg85]. If energy, i.e., the selfadjoint generator of time translations, is positive (or bounded from below), sharp localization described by projection-valued measures gives rise to causality problems. In modern terms, and especially in the formulation later clarified by Castrigiano, these problems concern the incompatibility of sharp localization with a natural requirement of causal (i.e. non-superluminal) propagation for localization probabilities. This strongly suggests that, in relativistic quantum theory, sharp notions of localization should be abandoned in favour of unsharp ones described by POVMs. In this case, effects, that is, positive operators bounded by the identity, $0\leq A(\Delta)\leq I$, rather than orthogonal projectors, are still associated with arbitrarily small regions $\Delta$ in the rest space $\Sigma$ of an observer, but the probability of finding the particle in a given bounded region is strictly $tr(\rho A(\Delta))<1$ for every state $\rho$, even though sharp states can be approximated arbitrarily well.
Recent works [Cas17, Mor23, Cas24, DRMo24, Cas25, CDM26] have shown that this strategy is in fact viable. A broad family of relativistic spatial localization observables has been constructed and analyzed for various elementary systems, with full compatibility with Poincar’e covariance and with natural causal requirements, especially in the unsharp setting when the requirement of positive energy is assumed. These developments indicate that Hegerfeldt’s issue does not rule out unsharp localization under the requirement of energy positivity, but rather rules out an excessively rigid, projection-based notion of localization.
@@
Actually, it is not even supposed that a whole unitary representation of the Poincar’e group exists. The relativistic perspective comes into play only to physically justify commutativity of localization effects associated with ”separated” regions, when interpreted as causally separated regions in the relativistic sense.
@@
The crucial point is that the commutativity requirement entering the Halvorson-Clifton theorem is usually regarded as self-evident from the perspective of local quantum physics, especially within the Araki-Haag-Kastler framework [Haa96, Ara09], where locality is encoded precisely by commutativity of observables associated with causally separated spacetime regions. If the spatial region $\Delta$ of a rest space $\Sigma$ where a particle is found at a certain time $t$ belongs to the spacetime region $\cal O$} where the localization-observable $A(\Delta)$ is localized (in AHK sense), then the commutativity assumption would indeed appear unavoidable. From this standpoint, Halvorson-Clifton’s result would represent an insurmountable obstruction to any notion of position of a quantum system at a given time in a reference frame. However, this line of reasoning already presupposes a specific mathematical representation of locality. It is therefore natural to investigate the interplay of commutativity and causal separation of the localization region, independently of the assumption that localization observables $A(\Delta)$ belong to AHK local algebras ${{\cal A}}({\cal O})$, in the strict sense, attached to arbitrarily small spacetime neighborhoods $\cal O$ of the corresponding spatial regions: ${\cal O}\supset\Delta$. A first step in this direction was taken by Busch [Bus99], building on a technical result with Singh [BuSi98], by analyzing the relation between commutativity and locality directly within the general theory of quantum measurement. In Busch’s approach an observable is localized in a region $\cal O$} of spacetime if it can be measured with instruments which work therein. This is also in the spirit of AHK structure, but in contrast to that framework, no further requirement about the structure (*, $C^{*}$, von Neumann algebra) is assumed for the family of observables localized in a region of spacetime. Similarly, no a priori requirements for families of observables associated with distinct regions of spacetime are supposed. In that setting, locality is formulated through operational principles such as the no-signaling condition and, in a stronger form, what is often called relativistic consistency [Bec20], namely invariance of the statistics of suitable consecutive measurements under interchange of their temporal order when the measurements take place in causally separated regions. For broad classes of measurement procedures – in particular for Lüders measurements and related efficient schemes – these principles imply commutativity of the relevant effects. This analysis is mathematically precise and conceptually important because it does not assume locality in the algebraic form of microcausality (i.e. commutativity), but rather attempts to derive commutativity from independent operational requirements.
@@
The above negative answer is not due to a failure of no-signaling or relativistic consistency, but rather to the physical meaning of localization for a quantum system which we define to be a particle. Let us assume that a localization observable is meant to describe a system with the propensity to be found in a unique place when an ideal position measurement is performed on a whole rest space. Then the elementary observable associated with detection in a region $\Delta$ is not naturally localized in an arbitrarily small spacetime neighborhood of $\Delta$. On the contrary, once one adopts the above stated local detectability principle, as we shall prove in Section 4.2, one is led to the conclusion that the elementary observable $A(\Delta),I-A(\Delta)$} is localized, in Busch’s sense, in a region that must contain the whole rest space where the global localization experiment is performed. As a consequence, two elementary localization observables associated with spatially disjoint regions are not localized in causally separated spacetime regions, and therefore the no-signaling and relativistic-consistency arguments leading to commutativity cannot even be triggered. In this sense, the commutativity requirement for localization effects does not follow from those more elementary locality principles, since the very hypotheses for employing these principles are not satisfied. This conclusion is physically tied to the particle character of the systems under consideration. A particle is assumed not to localize simultaneously in several (causally) disjoint regions in an ideal exhaustive detection experiment, but rather in a unique place. If one measures whether the particle is found in a region $\Delta$ within a complete position measurement on a rest space $\Sigma$, then one automatically obtains information about the complementary region $\Sigma\setminus\Delta$. From this viewpoint, the associated elementary observable is not confined to $\Delta$ alone: it carries information about the global detection arrangement on $\Sigma$. The failure of the commutativity argument is therefore not an exotic mathematical pathology, but rather a direct consequence of the physical meaning attributed to localization of a particle. If we do not accept this, we cannot accept the very notion of a particle.
These arguments provide a motivation for asserting that localization effects, for particles, do not belong to local algebras in the sense of AHK formulation, in the case where they are parts of a POVM which fills the whole rest space of a reference frame (a Cauchy surface).
@@
We introduce a notion of conditional localization associated with a bounded spatial region $\Delta_{0}$, interpreted as the spatial part of a laboratory. The idea is to construct, starting from a general localization observable defined on a full rest space, a new POVM normalized on $\Delta_{0}$ which describes the conditional probability of finding the system in a subregion $\Delta\subset\Delta_{0}$, given that the system has been detected in $\Delta_{0}$. The mathematical possibility of such construction relies upon a known result in quantum information theory called gentle measurement lemma. It permits one to prove that a specific type of POVMs we introduce in a laboratory normalized on bounded spatial region $\Delta_{0}$, measure the conditional probability of finding a particle in $\Delta\subset\Delta_{0}$ if we know that the system was detected in $\Delta_{0}$. This is all the more true the closer the probability of finding the particle in $\Delta_{0}$ is to $1$. These conditional POVMs are intrinsically local to the laboratory and, unlike the original localization observables on the whole rest space, are not affected by the same problem that prevents the use of no-signaling or relativistic-consistency arguments. We shall argue that pairs of such conditional POVMs associated with causally separated laboratories may, in principle, satisfy commutativity and belong to local algebras concentrated around the laboratory region in the AHK sense. This suggests that conditional localization may provide a mathematically and physically appropriate notion of local position measurement in relativistic quantum theory. Explicit examples of these local observables satisfying a commutativity requirement and constructed in terms of local energies will be presented is a forthcoming paper within the approach of local QFT.